how do i find the product of 0.06, -0.4 and -1.5
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

(0.06)(-0.4)(-1.5)
= (0.06)(-6)
= -0.36

Answer by ikleyn(52778)

.

            The solutions by   @CubeyThePenguin  is INCORRECT.

            I came to bring you the correct solution.


0.06*(-0.4)(-1.5) =

= 0.06*0.6 =

= 0.036.


Oho  (!)   "Tutor"   CubeyThePinguin   can not multiply numbers correctly  (!)


It is a news for me   (!)   How he  (or she)  is going to teach students  ?